# PhysioWave: A Multi-Scale Wavelet-Transformer for Physiological Signal Representation

### v2

@@ A.1 Self-supervised pre-training for Time Series
Self-supervised learning for time-series data has seen considerable advances with the development of models like MOMENT and OTIS, which are designed to handle the unique characteristics of time-series signals, such as non-stationarity, noise, and high variability across domains. Both models leverage large-scale pretraining to learn generalized representations for a wide range of tasks.
@@
The MOMENT model, for instance, introduces a self-supervised learning framework for time-series data using masked representation learning. It captures both local and global temporal dynamics through its ability to reconstruct masked input sequences. MOMENT has been shown to perform well in tasks like classification, anomaly detection, and forecasting, even without task-specific supervision. Additionally, it applies contrastive learning to model dependencies and trends in the data, addressing challenges such as long-range dependencies in time series sequences [2].
@@
OTiS, on the other hand, extends the idea of self-supervised learning to multi-domain time-series data. Its pre-training paradigm incorporates domain-specific tokenization and a dual masking strategy to handle the heterogeneity across time series from different domains, including EEG, audio, and financial data. By using a domain-specific tokeniser with learnable signatures, OTiS can capture domain-specific data characteristics while learning generalized features. This allows it to excel in various downstream tasks, such as classification, regression, and forecasting, across multiple domains [38].
@@
Both models highlight the importance of large-scale pretraining on diverse datasets to unlock foundational model capabilities. MOMENT and OTiS aim to generalize across domains, mitigating issues caused by domain-specific differences like sampling frequencies and inter-variate relationships. These models represent significant strides in time-series analysis, particularly in scenarios where labeled data is scarce but large, diverse, and unlabeled datasets are available for pretraining.
@@
A.2 Self-supervised pre-training for EEG signal
@@
Self-supervised learning has gained significant traction, particularly in natural language processing (NLP) and computer vision (CV), and has shown promising results in the domain of electroencephalography (EEG) [41]. In EEG, self-supervised learning models are leveraging the ability to pre-train models on large, unlabeled datasets, which can then be fine-tuned for specific downstream tasks. Two notable models in this domain are EEGPT and LaBraM, both of which advance self-supervised learning techniques for EEG signal analysis.
@@
EEGPT (EEG Pretrained Transformer) employs a masked autoencoder-based self-supervised learning strategy specifically designed for EEG data. This model is trained to predict masked segments of the EEG signal, enabling it to learn robust representations of the signal’s spatio-temporal features. The dual self-supervised learning approach in EEGPT combines masked reconstruction and spatio-temporal alignment, effectively addressing the challenges posed by low signal-to-noise ratio (SNR) and inter-subject variability in EEG. This method allows EEGPT to perform well across a variety of EEG-based tasks, particularly in the context of brain-computer interface (BCI) applications where accurate feature extraction is critical [5].
@@
LaBraM (Large Brain Model) is another significant model in self-supervised EEG pre-training. LaBraM focuses on learning universal EEG representations through unsupervised pre-training on large-scale EEG datasets from various domains. To address the challenge of cross-dataset learning, LaBraM segments EEG signals into channel patches and uses a vector-quantized neural spectrum prediction model to generate a compact neural tokenizer. This allows the model to learn generalized representations that can be fine-tuned on specific tasks, such as abnormality detection, emotion recognition, and gait prediction. The model’s ability to handle diverse EEG data with varying channel configurations and recording settings has been validated across multiple EEG-based tasks, demonstrating its robustness and scalability [4].
@@
Both models underscore the importance of large-scale pretraining in the EEG domain. By learning from vast amounts of unlabeled data, these self-supervised models can capture the rich, multi-scale features of EEG signals, which are crucial for a wide range of applications in health monitoring and BCI systems. The ability of EEGPT and LaBraM to generalize across datasets and tasks highlights the potential of self-supervised learning to transform the analysis of EEG signals, making deep learning techniques more applicable and effective in the field of bio-signal processing.
A.3 Self-supervised Pre-training for ECG Signal
@@
Self-supervised learning (SSL) has become a pivotal method in ECG foundation model development, addressing challenges such as the scarcity of labeled data and the need for robust generalization across diverse datasets. Recent advancements have introduced several effective SSL-based pre-training strategies for ECG data.
@@
One such approach is OpenECG, which created a large ECG benchmark with over 1.2 million 12-lead ECG recordings from 9 centers. The study evaluated three prominent SSL methods—SimCLR, BYOL, and MAE—using ResNet-50 and Vision Transformer (ViT) architectures. The results showed that pre-training on diverse datasets significantly improved model generalization, with BYOL and MAE outperforming SimCLR. These methods demonstrated superior effectiveness in learning feature-consistency and generative representations compared to contrastive methods, which require large datasets to perform optimally [42].
@@
Another notable model is ECG-Chat, a large ECG-language model designed for cross-modal cardiac diagnosis. It combines ECG waveform data with text reports using contrastive learning to align ECG features with medical text. The model was trained on a dataset that integrates both diagnosis and conversation tasks, achieving state-of-the-art results in ECG medical report generation. ECG-Chat also incorporates a novel method for mitigating hallucinations during report generation by integrating external cardiology knowledge via GraphRAG and DSPy components. This ensures that the generated ECG reports are grounded in clinical knowledge, enhancing accuracy and reliability [35].
@@
These models highlight the potential of self-supervised learning for enhancing ECG analysis and improving diagnostic tools. By leveraging large-scale ECG data and pre-training strategies, both OpenECG and ECG-Chat advance the capabilities of ECG models, making them more robust and adaptable across a range of clinical and research applications. They represent a significant step toward improving the accessibility and accuracy of AI-driven cardiovascular diagnostics.
Appendix B HYPERPARAMETER SETTINGS
@@
We employ AdamW with a weight decay of $0.01$ and moment coefficients $\beta_{1}=0.9$ and $\beta_{2}=0.98$. The learning rate is linearly warmed up from $5\times 10^{-7}$ to $5\times 10^{-5}$ over the first ten epochs and then follows a cosine decay to a floor of $1\times 10^{-6}$. Pretraining lasts for 50 epochs with a global batch size of 64 on 16 NVIDIA A100 GPUs, whereas all downstream experiments are carried out on 4 A100 GPUs. During downstream training, the same AdamW optimizer and the cosine scheduler is retained, but the learning rate is reduced to 1/10 of its pretraining value.
@@
Wavelet bases {sym4, sym5, db6, coif3, bior4.4}
@@
Decoder
Decoder embed dim 256
Decoder depth / heads 8 / 8
@@
Decoder
Decoder embed dim 256
Decoder depth / heads 8 / 8
@@
Appendix C PRETRAINING DATASETS DESCRIPTION
C.1 Pretraining datasets for PhysioWave-ecg
In the pretraining of PhysioWave-ecg, we process ECG data to ensure it is suitable for model training. The process involves several key steps, including filtering, windowing, and handling different sampling rates.
First, we apply a bandpass filter (0.5–40 Hz) to remove unwanted low-frequency noise and high-frequency artifacts. Additionally, a 50 Hz notch filter is applied to eliminate power line interference. These filtering steps are crucial for improving the quality of the ECG signals.
To handle the data, we use a sliding window technique. The ECG signals, with 12 leads and a sampling rate of 500 Hz, are divided into overlapping windows of a fixed size (e.g., 1024 samples). If the signal length is insufficient for a full window, we apply zero-padding to maintain consistency across the dataset.
For ECG recordings with a sampling rate lower than 500 Hz, we upsample the signal to 500 Hz to ensure uniformity across the dataset. This is important for maintaining the integrity of the signal while ensuring that all data points are processed at the same sampling frequency.
@@ C.1 Pretraining datasets for PhysioWave-ecg
C.2 Pretraining datasets for PhysioWave-emg
@@
For the pretraining of the PhysioWave-emg model, we process the raw EMG data through several critical preprocessing steps to ensure its suitability for training. First, a bandpass filter is applied to remove unwanted noise and artifacts, retaining frequencies between 20 Hz and 450 Hz, which are relevant for EMG analysis. A notch filter is then used to suppress 50 Hz power line interference, a common source of noise in EMG recordings. After filtering, we normalize the signals using z-score normalization, ensuring that each EMG channel has a mean of zero and a standard deviation of one, standardizing the input scale for machine learning models.
To prepare the data for training, we apply a sliding window approach, dividing the EMG signal into overlapping windows of 1024 samples with a step size of 512. If the sampling rate of the raw signal is below 2000 Hz, the data is upsampled to 2000 Hz to ensure consistency. Additionally, if the number of channels is less than 16, the data is padded with zeros to match the required dimensionality. Finally, the processed data is saved in an HDF5 format, allowing for efficient storage and incremental updates. This ensures easy access to the data and facilitates the addition of new data without overwriting existing information.
@@
Appendix D Downstream Datasets Description
@@
D.1 Downstream datasets for PhysioWave-ecg
@@
D.2 Downstream datasets for PhysioWave-emg
@@
D.3 Downstream datasets for multi-modal tasks
@@
Appendix E VISUALIZATION
E.1 Visualization of sEMG Pretraining

### v4

@@ B.1 Self-supervised pre-training for Time Series
Self-supervised learning for time series data has seen considerable advances with the development of models such as MOMENT and OTIS, which are designed to handle the unique characteristics of time-series signals, such as non-stationarity, noise, and high variability across domains. Both models leverage large-scale pretraining to learn generalized representations for a wide range of tasks.
@@
The MOMENT model, for instance, introduces a self-supervised learning framework for time series data using masked representation learning. Capture both local and global temporal dynamics through its ability to reconstruct masked input sequences. MOMENT has been shown to perform well in tasks like classification, anomaly detection, and forecasting, even without task-specific supervision. Furthermore, it applies contrastive learning to model dependencies and trends in the data, addressing challenges such as long-range dependencies in time series sequences [2].
@@
OTiS, on the other hand, extends the idea of self-supervised learning to multi-domain time-series data. Its pre-training paradigm incorporates domain-specific tokenization and a dual masking strategy to handle the heterogeneity across time series from different domains, including EEG, audio, and financial data. Using a domain-specific tokeniser with learnable signatures, OTiS can capture domain-specific data characteristics while learning generalized features. This allows it to excel in various downstream tasks, such as classification, regression, and forecasting, in multiple domains [36].
@@
Both models highlight the importance of large-scale pretraining on diverse datasets to unlock basic modeling capabilities. MOMENT and OTiS aim to generalize across domains, mitigating issues caused by domain-specific differences such as sampling frequencies and intervariate relationships. These models represent significant strides in time-series analysis, particularly in scenarios where labeled data are scarce, but large, diverse, and unlabeled datasets are available for pre-training.
@@
B.2 Self-supervised pre-training for EEG signal
@@
Self-supervised learning has gained significant traction, particularly in natural language processing (NLP) and computer vision (CV), and has shown promising results in the domain of electroencephalography (EEG) [39]. In EEG, self-supervised learning models are leveraging the ability to pretrain models on large, unlabeled datasets, which can then be fine-tuned for specific downstream tasks. Two notable models in this domain are EEGPT and LaBraM, both of which advance self-supervised learning techniques for the analysis of EEG signals.
@@
EEGPT (EEG Pretrained Transformer) employs a masked autoencoder-based self-supervised learning strategy specifically designed for EEG data. This model is trained to predict masked segments of the EEG signal, enabling it to learn robust representations of the signal’s spatiotemporal features. The dual self-supervised learning approach in EEGPT combines masked reconstruction and spatiotemporal alignment, effectively addressing the challenges posed by low signal-to-noise ratio (SNR) and intersubject variability in EEG. This method allows EEGPT to perform well in a variety of EEG-based tasks, particularly in the context of brain-computer interface (BCI) applications where accurate feature extraction is critical [5].
@@
LaBraM (Large Brain Model) is another significant model in self-supervised EEG pre-training. LaBraM focuses on learning universal EEG representations through unsupervised pre-training on large-scale EEG datasets from various domains. To address the challenge of cross-dataset learning, LaBraM segments EEG signals into channel patches and uses a vector-quantized neural spectrum prediction model to generate a compact neural tokenizer. This allows the model to learn generalized representations that can be fine-tuned on specific tasks, such as abnormality detection, emotion recognition, and gait prediction. The model’s ability to handle diverse EEG data with varying channel configurations and recording settings has been validated in multiple EEG-based tasks, demonstrating its robustness and scalability [4].
@@
Both models underscore the importance of large-scale pretraining in the EEG domain. By learning from vast amounts of unlabeled data, these self-supervised models can capture the rich, multi-scale features of EEG signals, which are crucial for a wide range of applications in health monitoring and BCI systems. The ability of EEGPT and LaBraM to generalize across datasets and tasks highlights the potential of self-supervised learning to transform the analysis of EEG signals, making deep learning techniques more applicable and effective in the field of biosignal processing.
B.3 Self-supervised Pre-training for ECG Signal
@@
Self-supervised learning (SSL) has become a key method in the development of foundational ECG models, addressing challenges such as the scarcity of labeled data and the need for robust generalization in diverse datasets. Recent advancements have introduced several effective SSL-based pretraining strategies for ECG data.
@@
One such approach is OpenECG, which created a large ECG benchmark with over 1.2 million 12-lead ECG recordings from 9 centers. The study evaluated three prominent SSL methods: SimCLR, BYOL, and MAE using ResNet-50 and Vision Transformer (ViT) architectures. The results showed that pretraining on diverse datasets significantly improved model generalization, with BYOL and MAE outperforming SimCLR. These methods demonstrated superior effectiveness in learning feature consistency and generative representations compared to contrastive methods, which require large datasets to perform optimally [44].
@@
Another notable model is ECG-Chat, a large ECG-language model designed for cross-modal cardiac diagnosis. It combines ECG waveform data with text reports using contrastive learning to align ECG features with medical text. The model was trained on a data set that integrates both diagnosis and conversation tasks, achieving state-of-the-art results in the generation of medical reports from ECG. ECG-Chat also incorporates a novel method to mitigate hallucinations during report generation by integrating external cardiology knowledge via GraphRAG and DSPy components. This ensures that the generated ECG reports are grounded in clinical knowledge, enhancing accuracy and reliability [35].
@@
These models highlight the potential of self-supervised learning to improve ECG analysis and improve diagnostic tools. Using large-scale ECG data and pretraining strategies, both OpenECG and ECG-Chat advance the capabilities of ECG models, making them more robust and adaptable across a range of clinical and research applications. They represent a significant step toward improving the accessibility and accuracy of AI-driven cardiovascular diagnostics.
Appendix C HYPERPARAMETER SETTINGS
@@
We employ AdamW optimizer with a weight decay of $0.01$ and moment coefficients $\beta_{1}=0.9$ and $\beta_{2}=0.98$. The learning rate is linearly warmed up from $5\times 10^{-7}$ to $5\times 10^{-5}$ over the first ten epochs and then follows a cosine decay to a floor of $1\times 10^{-6}$. Pretraining lasts for 50 epochs with a global batch size of 64 on 16 NVIDIA A100 GPUs, whereas all downstream experiments are carried out on 4 A100 GPUs. During downstream training, the same AdamW optimizer and cosine scheduler are retained, but the learning rate is reduced to 1/10 of its pretraining value.
@@
Wavelet bases {sym4, db4, db6, coif2}
@@
Reconstruction Head
Hidden dimensions [256]
Output dimension (patch dim) 768 (12 channels $\times$ 64 timesteps)
Activation function GELU
Dropout rate 0.1
Layer normalization True
@@
Reconstruction Head
Hidden dimensions [256]
Output dimension (patch dim) 512 (8 channels $\times$ 64 timesteps)
Activation function GELU
Dropout rate 0.1
Layer normalization True
@@
Appendix D PRETRAINING DATASETS DESCRIPTION
D.1 Pretraining datasets for PhysioWave-ecg
In the pretraining of PhysioWave-ECG, we streamline the data processing pipeline by leveraging the inherent filtering capabilities of our wavelet-based architecture. Unlike conventional approaches that require extensive preprocessing with bandpass filters, notch filters, and baseline wander removal, our adaptive wavelet decomposition naturally handles these signal conditioning tasks within the model itself. The process involves two main steps. First, we use a sliding window technique where ECG signals with 12 leads are divided into overlapping windows of 2048 samples. If the signal length is insufficient for a full window, we apply zero-padding to maintain consistency across the dataset. Second, for ECG recordings with sampling rates different from our target 500 Hz, we resample the signals to ensure uniformity across the dataset. Notably, we intentionally omit traditional filtering operations (such as 0.05-100 Hz bandpass filtering or 60 Hz notch filtering) from our preprocessing pipeline. This is because our multi-level wavelet decomposition with learnable basis functions serves as an adaptive, data-driven filtering mechanism. The wavelet transform inherently provides frequency-selective decomposition, effectively separating signal components across different frequency bands. Our learnable wavelet kernels adapt during training to optimally filter noise, remove baseline wander, and preserve clinically relevant signal features—all without manual filter design or fixed frequency cutoffs. This approach not only simplifies the preprocessing pipeline but also allows the model to learn task-specific filtering strategies that may be more effective than generic, hand-crafted filters.
@@ D.1 Pretraining datasets for PhysioWave-ecg
D.2 Pretraining datasets for PhysioWave-emg
@@
For the pretraining of the PhysioWave-EMG model, we adopt a streamlined preprocessing approach that leverages our wavelet-based architecture’s inherent signal processing capabilities. Unlike traditional EMG processing pipelines that require extensive filtering operations, our adaptive wavelet decomposition naturally handles noise suppression and feature extraction within the model itself.Our preprocessing focuses on data standardization rather than filtering. We apply z-score normalization to ensure each EMG channel has a mean of zero and a standard deviation of one, standardizing the input scale.
Notably, we intentionally omit conventional bandpass filtering (20-450 Hz) and notch filtering (50/60 Hz power line interference) from our preprocessing. This is because our learnable wavelet transform provides adaptive, data-driven filtering that automatically separates relevant EMG signal components from noise and artifacts. The multi-level wavelet decomposition with trainable kernels learns to isolate muscle activation patterns while suppressing interference, effectively replacing manual filter design with learned, task-specific signal conditioning.To prepare the data for training, we apply a sliding window approach, dividing the EMG signal into overlapping windows of 1024 samples with a step size of 512. For consistency across datasets, we standardize all signals to 2000 Hz sampling rate through resampling when necessary.Regarding channel configuration, we standardize all inputs to 8 channels to maintain architectural consistency and computational efficiency. For multi-channel EMG recordings that exceed 8 channels, we perform channel selection during preprocessing, retaining the 8 most informative channels based on signal quality metrics such as signal-to-noise ratio and muscle activation patterns. This selective approach ensures we capture the most relevant EMG information while maintaining a uniform input dimension that balances representational capacity with computational efficiency.Finally, the processed data is saved in HDF5 format, enabling efficient storage and incremental updates. This structured approach not only simplifies the preprocessing pipeline but also allows the wavelet-based model to adaptively learn optimal signal processing strategies directly from the data, potentially discovering more effective filtering and feature extraction methods than traditional fixed preprocessing techniques.
@@
Appendix E Downstream Datasets Description
@@
E.1 Downstream datasets for PhysioWave-ecg
@@
E.2 Downstream datasets for PhysioWave-emg
@@
E.3 Downstream datasets for multi-modal tasks
@@
Appendix F VISUALIZATION
F.1 Mathematical Properties Validation
We conducted a comprehensive mathematical analysis of the learned wavelet kernels extracted from our PhysioWave model, examining 10 filter pairs (5 wavelet bases × 2 filters) across multiple decomposition levels. Our analysis reveals that the model has successfully developed task-adaptive signal decomposition kernels that prioritize empirical performance over strict mathematical constraints.
Figure 6: Learned wavelet kernel shapes from PhysioWave model. The upper panel shows all 10 learned filters (5 low-pass and 5 high-pass) with vertical offsets for clarity. Lower panels display individual filter characteristics, demonstrating smooth low-pass filters for signal approximation and oscillatory high-pass filters for detail extraction.
F.1.1 Key Mathematical Properties
Energy Conservation: As shown in Figure 6, all learned filters demonstrate perfect unit energy normalization (100% of filters with $w||_{2}=1.0$), ensuring stable signal decomposition and preventing amplitude drift during multi-level analysis. This property is crucial to maintain the integrity of the signal at all decomposition levels[59].
Compact Support: The filters exhibit excellent spatial localization with average support lengths of 15.6 samples within the 16-sample window (97.5% utilization), demonstrating efficient use of the receptive field without boundary artifacts. This compact support enables precise temporal localization of EMG events[60].
Figure 7: Frequency domain analysis of learned wavelets. (a) Magnitude spectrum showing clear frequency separation between low-pass and high-pass filters. (b) Phase spectrum demonstrating phase coherence. (c) Energy distribution heatmap across frequency bins. (d) Frequency characteristics scatter plot showing clustering of filter responses in physiologically relevant bands.
Frequency Selectivity: Clear frequency differentiation emerged through training (Figure 7):
  • Low-pass filters: Frequency centroids at 6.0–7.1 (normalized) with broader bandwidths (5.6–6.4), optimized for capturing EMG envelopes and removing high-frequency noise.
  • High-pass filters: Higher frequency centroids at 7.7–7.8 with narrower bandwidths (3.0–5.3), specialized for detecting rapid muscle activation transients.
F.1.2 Task-Specific Adaptations
Our analysis reveals that the learned kernels have evolved beyond traditional wavelet constraints to become physiologically-informed decomposition filters. While classical wavelets maintain strict mathematical properties (zero mean, vanishing moments), our filters demonstrate data-driven optimization that better captures EMG signal characteristics:
Figure 8: Comparison between learned filters and standard wavelets. Each subplot shows a learned filter (blue solid line) overlaid with its most similar standard wavelet (red dashed line), with correlation coefficients (r) indicating the degree of similarity. The low correlation values (r = 0.033–0.104) demonstrate that our model has discovered novel, task-specific decomposition kernels distinct from traditional wavelet families.
  1. 1.
    Adaptive Mean Values: Non-zero mean values (0.017–0.149) allow the filters to better track EMG baseline shifts and muscle fatigue patterns, which often present as DC components in real physiological recordings.
  2. 2.
    Specialized Frequency Response: The concentration of high-pass filter responses around 7.7–7.8 (normalized frequency) corresponds to the dominant frequency band of motor unit action potentials in EMG signals (typically 20–150 Hz), demonstrating automatic discovery of physiologically relevant frequency bands.
  3. 3.
    Heterogeneous Smoothness: Variable smoothness characteristics (1st derivative std: 0.104–0.465) reflect adaptation to the non-stationary nature of EMG signals, with smoother filters for baseline tracking and sharper filters for spike detection.
F.1.3 Comparison with Standard Wavelets
As illustrated in Figure 8, the learned filters show minimal correlation (r = 0.033–0.104) with standard wavelet families (db4, db6, sym4, sym5, coif3). This divergence is not a limitation but rather evidence of successful adaptation to the unique characteristics of physiological signals. Traditional wavelets were designed for general signal processing, whereas our learned filters have automatically discovered decomposition patterns specifically optimized for EMG signal structures.
F.1.4 Empirical Validation
The departure from traditional wavelet properties represents a principled trade-off between mathematical elegance and empirical performance. Our learned filters achieve:
  • Superior classification accuracy (94.2% on EMG gesture recognition)
  • Enhanced noise robustness in clinical settings
  • Better generalization across subjects compared to fixed wavelet bases
This shows that for biosignal processing applications, task-optimized learned filters outperform mathematically constrained traditional wavelets, validating our approach of allowing the model to discover optimal decomposition kernels through end-to-end learning. The filters have effectively learned to extract the most discriminative features from EMG signals while maintaining the multiresolution analysis framework essential for physiological signal processing.
F.2 Visualization of sEMG Pretraining